OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (shamim):

required equation is y=mx+c

OpenStudy (anonymous):

ya

OpenStudy (shamim):

here slope m=-1

OpenStudy (shamim):

x=-1 y=4

OpenStudy (anonymous):

so what would c be

OpenStudy (shamim):

4=-1(-1)+c

OpenStudy (anonymous):

so c can be either 4 or 6 right

OpenStudy (shamim):

no

OpenStudy (anonymous):

im confused

OpenStudy (shamim):

4=1+c c=4-1 c=3

OpenStudy (anonymous):

so now i plug it in right

OpenStudy (shamim):

-1(-1)=1

OpenStudy (shamim):

ya

OpenStudy (anonymous):

i got x – 4y = –1

OpenStudy (shamim):

y=-1x+3 y=-x+3 x+y=3

OpenStudy (shamim):

did u get it